# Reborn: Reinforcement-Learned Boundary Segmentation with Iterative Training for Unsupervised ASR

### v1

@@ 5.3.1   Reward Design and Behavior Cloning
Table 5: We compare different reward functions and the effect of BC initialization with first iteration results on LibriSpeech. Row (c) is Reborn.
Ablation Stage 1 PER (%) Stage 2 PER (%) Silence Ratio dev-clean test-clean dev-clean test-clean (a). $R_{\text{ppl}}$ 14.0 14.7 12.2 12.9 0.13 (b). $R_{\text{ppl}}+R_{\text{edit}}$ 13.3 13.8 11.7 12.1 0.07 (c). $R_{\text{ppl}}+R_{\text{edit}}+R_{\text{len}}$ 12.7 12.9 11.7 11.9 0.05 (d). Reborn w/o BC 14.8 14.9 14.1 14.2 0.05
We verify the effectiveness of our three proposed reward functions and behavior cloning initialization in Table 5. While the perplexity difference reward alone is sufficient to guide the segmentation model, adding edit distance and length difference rewards improves performance further. As previously mentioned, only using the perplexity difference reward will make the segmentation model predict more segments that lead to phoneme prediction sequences with more $<$silence$>$ tokens, which have a lower PPL but do not correspond to better phoneme prediction results. To demonstrate this, we show the silence ratio in the de-duplicated phoneme prediction sequence per utterance. We find that the silence ratio when only using $R_{\text{ppl}}$ (row (a)) is 0.13, which is much higher than the silence ratio of ground truth phoneme transcription sequences, which is 0.07. Adding $R_{\text{edit}}$ (row (b)) and $R_{\text{len}}$ (row (c)) makes the silent ratio closer to the ground truth silence ratio.
We also find that removing BC (row (d)) worsens PER compared with using BC (row (c)). Additionally, BC greatly speeds up RL training convergence, as shown in Figure 4 (detailed in Appendix B).
@@ 5.3.2   Iterative Training
In Figure 3, we show the PER on LibriSpeech test-clean split during the iterative training process of Reborn. We observe that after training the segmentation model in the first iteration, the PER drastically drops by 6.4% compared with wav2vec-U, which is used as the initialization of the phoneme prediction model in iteration one. This shows that the quality of the segmental structure’s boundaries is important to the performance of UASR. Afterward, the PER of Reborn decreases steadily across iterative training, showing that training the segmentation model and the phoneme prediction model is critical to improving the performance of Reborn. We find that the PER does not drop significantly after the fifth iteration.
@@ 6 Conclusion
Impact Statements
@@
Unsupervised automatic speech recognition (UASR) - predicting the textual transcriptions for the given speech signals using only unpaired speech and text data - is highly attractive. The existence of thousands of low-resourced languages spoken globally is the major reason. This paper presents Reborn, a novel unsupervised ASR training algorithm. We foresee that Reborn has a great potential to make low-resource languages more accessible to train ASR systems, making speech technology more accessible and boundaryless.
However, our training method may reinforce the bias in the dataset. This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A wav2vec-U
@@
wav2vec-U (Baevski et al., 2021) is a popular UASR framework whose variants (Gao et al., 2023; Liu et al., 2022) have achieved SoTA performance on multiple UASR benchmarks. Following prior works (Liu et al., 2018), wav2vec-U also uses adversarial training to train the phoneme prediction model. The wav2vec-U framework takes the raw waveform as the input and predicts phoneme transcription via three stages: (1) the feature extraction stage, (2) the feature preprocessing stage, and (3) the phoneme prediction stage.
@@
Stage (1) extracts speech features using a self-supervised speech foundation model, wav2vec 2.0 (Baevski et al., 2020). For a spoken utterance, we denote the extracted feature as $X=[x_{1},x_{2},...,x_{T}]\in\mathbb{R}^{T\times d}$, where $T$ is the time dimension and $d$ is the feature dimension. Stage (2), the feature preprocessing stage, uses $k$-means clustering and additional heuristics, namely, adjacent pooling, to define the boundary of the segmental structures in the speech features. The features in the same segmental structure are mean-pooled into one feature embedding to reduce sequence length to match phoneme sequences better. The pooled features’ dimension is reduced using PCA. We denote the output from the feature preprocessing stage as $S=[s_{1},s_{2},...,s_{T^{\prime}}]\in\mathbb{R}^{T^{\prime}\times d^{\prime}}$. In Stage (3), the phoneme prediction model takes the pooled speech feature as input and predicts the phoneme sequence $Y=[y_{1},y_{2},...,y_{T^{\prime}}]$. Last, a de-duplication step removes the consecutive duplicate phoneme predictions from $Y$, resulting in a shortened phoneme sequence $Y^{\prime}=[y^{\prime}_{1},y^{\prime}_{2},...,y^{\prime}_{M}]$ without consecutive duplicate phonemes. $Y^{\prime}$ will be used as the input to fool the discriminator in adversarial training, which loss is defined as follows:
$\mathcal{L}_{GAN}=$ $\min_{G}\max_{D}\underset{Z\sim\mathcal{Z}}{\mathbb{E}}[\log(\mathcal{D}(Z))]+\underset{X\sim\mathcal{X}}{\mathbb{E}}[\log(1-\mathcal{D}(\mathcal{G}(X)))]$ (5)
@@
where $Z$ is a phoneme sequence sampled from the text corpora $\mathcal{Z}$ and $X$ is the speech signal from the speech corpora $\mathcal{X}$. Besides the min-max objective in the adversarial training, wav2vec-U (Baevski et al., 2021) also adopts the three regularized objectives for better convergence and performance. First, the gradient penalty $\mathcal{L}_{gp}$ (Gulrajani et al., 2017) is applied on the discriminator $\mathcal{D}$ to stabilize training, where the input of $\mathcal{D}$ is the linear combination of a generated sample $Y$ and a real sample $Z$.
$\mathcal{L}_{gp}=\underset{Y,Z,\alpha\sim U(0,1)}{\mathbb{E}}\left[\left(\left\|\nabla\mathcal{D}(\alpha Y+(1-\alpha)Z)\right\|_{2}-1\right)^{2}\right]$ (6)
@@
Next, the smoothness penalty $\mathcal{L}_{sp}$ is computed across consecutive segments of the generated sequence, promoting the generation of closely aligned neighboring outputs by the model.
$\mathcal{L}_{sp}=\sum_{(y_{t},y_{t+1})\in Y}\left\|y_{t+1}-y_{t}\right\|^{2}$ (7)
@@
Furthermore, the phoneme diversity loss $\mathcal{L}_{pd}$ is applied to encourage higher vocabulary usage of the generated sequences. As shown in Equation 8, the entropy of the averaged softmax distribution of the generator outputs ($H_{\mathcal{G}}(\mathcal{G}(S))$) is maximized over the phoneme vocabulary across a batch B.
@@
In this work, we utilize the same training objective from wav2vec-U to perform GAN training (§ 3.2)
Appendix B Impact of Behavior Cloning on Model Convergence
Figure 4: Training Dynamics of BC Pre-Trained vs. Scratch Models. This figure depicts the evolution of Phoneme Error Rate (PER) over training epochs, tested on the test-clean set, to showcase the enhanced convergence speed and efficiency achieved with BC pre-training.
In this ablation study, we rigorously examine the influence of Behavior Cloning (BC) pre-training on the convergence dynamics and overall efficacy of our segmenter model. Our empirical findings reveal a marked improvement in sample efficiency for models subjected to BC pre-training, which is a critical aspect in enhancing model performance. Notably, these models demonstrate a notably faster convergence rate as compared to their counterparts trained from scratch.
As depicted in Figure 4, the training trajectories of the models, with and without BC pre-training, are contrasted. This visualization distinctly underscores the difference in PER across various training epochs. The curves elucidate not only the expedited convergence but also the superior performance attainable with BC pre-trained models.
It is observed that segmenters solely pre-trained through BC do not exhibit a substantial improvement in PER initially. However, this pre-training imbues the model with a more refined capability for boundary prediction. In subsequent RL training phases, this enhanced boundary prediction proficiency translates into a swifter convergence trajectory, culminating in superior final performance metrics. This phenomenon highlights the nuanced yet significant role of BC pre-training in the context of unsupervised automatic speech recognition, underscoring its potential as a pivotal step in the model optimization process.
Appendix C Implementation Details
C.1 Segmentaiton model

### v2

@@ 5.3.1 Reward Design and Behavior Cloning
We verify the effectiveness of our three proposed reward functions and behavior cloning initialization in Table 5. While the perplexity difference reward alone is sufficient to guide the segmentation model, adding edit distance and length difference rewards improves performance further. As previously mentioned, using only the perplexity difference reward will make the segmentation model predict segments overly focusing on lowering PPL. And a phoneme sequence with a lower perplexity may not always be a better transcription. To deal with such overfitting problem, we design the two regularization rewards, namely, $R_{edit}$ and $R_{len}$, to make the model learn to segment for lower perplexity while grounded on a reasonable transcription from the previous iteration. Our results in Table 5 further evidence the effectiveness of the two rewards. Although having a higher perplexity, using the additional regularized rewards has successfully increased the performance with lower PER (row (c) vs. row (a)). We also find that removing BC (row (d)) worsens PER compared with using BC (row (c)). Additionally, BC greatly speeds up RL training convergence, as shown in Figure 3.
@@ 5.3.2 Iterative Training
In Figure 2, we show the PER on LibriSpeech test-clean split during the iterative training process of Reborn. We observe that after training the segmentation model in the first iteration, the PER drastically drops by 6.4% compared with wav2vec-U, which is used to initialize the phoneme prediction model in iteration one. This shows that the quality of the segmental structure’s boundaries is important to the performance of UASR. Afterward, the PER of Reborn decreases steadily across iterative training, showing that training the segmentation model and the phoneme prediction model is critical to improving the performance of Reborn. We find that the PER does not drop significantly after the fifth iteration. Additionally, we provide more evidence to show that the phoneme predictors are iteratively polished in Appendix C.4.
@@ 6 Conclusion
Appendix / supplemental material
Appendix A Limitations
Recent advancements in unsupervised ASR are still in the developmental stages, and we have not yet implemented Reborn in more realistic or noisy environments. Additionally, the iterative training paradigm may amplify any existing biases in the dataset, an aspect that remains unexplored in this study. Furthermore, if the phoneme prediction model is poorly initialized, Reborn may not be able to provide huge performance improvements. In this work, lexicons are required to perform phonemicization on the unpaired text. Consequently, learning the unsupervised ASR system directly from speech to word-level tokens remains an open problem.
Appendix B Broader Impacts
@@
Unsupervised automatic speech recognition (UASR) - predicting the textual transcriptions for the given speech signals using only unpaired speech and text data - is highly attractive. The existence of thousands of low-resourced languages spoken globally is the major reason. This paper presents Reborn, a novel unsupervised ASR training algorithm. We foresee that Reborn has great potential to make low-resource languages more accessible to train ASR systems, making speech technology more accessible and boundaryless.
Appendix C Additional Results
C.1 Impact of Behavior Cloning on Model Convergence
Figure 3: Comparison of training dynamics between BC pre-trained and from-scratch models. This figure shows the evolution of Phoneme Error Rate (PER) across training epochs on the test-clean split of LibriSpeech, illustrating the faster convergence and improved efficiency achieved with BC pre-training.
In this ablation study, we investigate the impact of Behavior Cloning (BC) pre-training on the performance of our segmenter model. Our results show that BC pre-training enhances sample efficiency and accelerates convergence compared to training from scratch. The training curves in Figure 3 highlight these differences, showing both faster convergence and improved performance with BC pre-training.
The segmenter solely pre-trained through BC exhibits noticeable improvements, but this is not yet substantial in terms of Phoneme Error Rate (PER) initially. However, this pre-training boosts the model’s boundary prediction capabilities, accelerating learning in later reinforcement learning phases. This indicates that BC pre-training is essential for enhancing our unsupervised automatic speech recognition framework, making it a crucial optimization step.
C.2 Reward Design Details
When training reinforcement learning, various methods have demonstrated that incorporating regularization terms can enhance model training stability [51, 52]. In response to this insight, we have integrated both edit-distance and length rewards as forms of regularization in our training regime. To identify the most effective reward configuration, we employed an unsupervised metric (Eq. 10) for selection. Detailed unsupervised metric scores and further evaluation results of our reward configurations are presented in Table 6.
Table 6: Ablation study results of various configurations on the unsupervised metric assessed using validation set and phoneme error rates evaluated on the LibriSpeech dataset.
Configuration Stage 1 Stage 2
Unsupervised Dev-clean Test-clean Unsupervised Dev-clean Test-clean
Metric (↓) PER (%) PER (%) Metric (↓) PER (%) PER (%)
$R_{\text{ppl}}$ 256621.6 14.0 14.7 206312.79 12.2 12.9
$R_{\text{ppl}}+R_{\text{edit}}$ 251476.6 13.3 13.8 198447.93 11.7 12.1
$R_{\text{ppl}}+R_{\text{edit}}+R_{\text{len}}$ 243419.4 12.7 12.9 191043.37 11.7 11.9
Reborn w/o BC 252302.3 14.8 14.9 237658.23 14.1 14.2
Our findings, as shown in Table 5, indicate that using perplexity (PPL) alone as a reward yields the lowest perplexity scores upon model convergence. However, this approach also results in higher scores on the unsupervised metric, implying a decline in model performance. Further analysis reveals that the PER increases when the PPL reward is the sole metric. This suggests that, without regularization, the model tends to minimize perplexity by producing more lengthy and less accurate transcriptions, which do not align well with the expected outputs.
Table 7: The table for stability analysis. Each method is trained for 5 runs on LibriSpeech and evaluated on the test-clean split. Reborn performs steadily well while being fully converged in both of the stages. The ablations show the effectiveness of applying boundary merging and parameter initialization from the previous iteration for the stage 2 GAN training. The abbreviation "adj." indicates adjacent pooling, which is the 2nd stage of pooling in wav2vec-U (see Appendix D). The "Freq." in the last column is the number of boundaries per second.
Ablation Type mean PER $\pm$ std (%) %-converged ($\uparrow$) Freq. (Hz) Initial stage wav2vec-U ($k$-means-only) GAN $>100$ 0% 28.5 wav2vec-U ($k$-means + adj.) GAN 20.1 $\pm$ 0.6 100% 14.3 Iteration 1 Reborn stage 1 RL 12.9 $\pm$ 0.7 100% 16.3 Reborn stage 2 GAN 12.0 $\pm$ 0.1 100% 11.6          w/o boundary merging GAN 13.9 $\pm$ 2.1 100% 16.3          w/ random initialization GAN 14.1 $\pm$ 0.4 100% 11.6 Topline Train from oracle boundary GAN 6.64 $\pm$ 0.2 100% 11.9
C.3 Stability Analysis
In unsupervised ASR, stability has always been one of the major concerns since the unlabeled scenario makes the problem difficult and the training unstable. In Table 7, we demonstrate that Reborn is stable by calculating the mean PER and the corresponding standard deviations, as well as the converged rate. We introduce that a converged experimental run is categorized by its result yielding PER $<40\%$, following Baevski et al. [4]. As represented in the table, Reborn generates significantly better performance while being stable with a 100% converged rate in both stages.
Next, we focus on the ablations of Reborn stage 2 GAN training in Table 7. When the boundary merging is not applied, the GAN training on the original Reborn-segmented features is still better than the initial stage. However, the average performance and the standard deviation are larger. The observation indicates that the training is slightly unstable, but does not diverge too much since it is still fully converged. On the other hand, when we train the next-iteration GAN with completely random initialized parameters, the performance degrades even if we train the model for more steps. These two ablations directly demonstrate the effectiveness of adopting boundary merging and parameter initialization from the previous iteration, while the GAN training can still work on the original boundary or with randomly initialized parameters.
Furthermore, we find out that the instability might be highly correlated to the boundary frequency (the last column in Table 7), and this observation also corresponds with the prior works [4, 28]. Just as we performed in Table 1, we show the topline result of training GAN from the oracle boundary in the last row. We contribute the good result to the disappearance of the length mismatch issue between the segmented feature and its textual transcription. In the Initail stage, when we use only the $k$-means for segmentation instead of the two-stage $k$-means + adjacent segmentation (or "$k$-means-based segmentation" in the main text) in wav2vec-U, the length mismatch issue arises and we can not obtain reasonable results.
As for Reborn, the boundary without merging has a higher frequency compared with the wav2vec-U segmentation. However, using the original boundary for GAN training can still yield fully converged results with lower PER. We contribute this result to the improved quality of Reborn boundary, showing the importance and effectiveness of our segmentation learning via RL.
C.4 Iteratively Polished Phoneme Predictors
Table 8: We demonstrate that the Reborn phoneme predictors are gradually improved though our iterative training. Each of the phoneme predictors takes the same oracle-segmented features as input.
Phoneme Predictor PER (%) Relative
(evaluate on oracle) Gain ($\uparrow$)
Initial stage
wav2vec-U 13.8 0%
Reborn
Iteration 1 10.9 21%
Iteration 2 9.9 28%
Iteration 3 7.6 45%
Iteration 4 7.4 46%
Iteration 5 8.7 37%
Topline
Train from oracle 6.4 54%
In Reborn, we introduce an iterative paradigm that can gradually polish the segmentation model and the phoneme predictor in turn. In Table 8, we wish to provide more evidence to directly demonstrate that the phoneme predictors are actually improved through the iterations. We carry this out by feeding the identical oracle-segmented features to the phoneme predictors in different stages. Given the same ideally-segmented features, a better phoneme predictor should yield a better performance as well. In Table 8, we show that the phoneme predictor in the initial stage gives 13.8% PER when evaluating on the oracle-segmented features. Reborn’s iterative learning paradigm gradually improves the phoneme predictor, and the performance finally achieves the best in the 4th iteration by 7.4% PER. The relative performance gain compared with the initial stage is over 45%, as presented in the table. The result is even 1% close to the topline listed in the bottom row, indicating the effectiveness of Reborn’s iterative training.
C.5 Generalizability across Different Speech Foundation Models
Table 9: We implement Reborn across different speech foundation models on LibriSpeech. The results are evaluated on test-clean. Reborn has exhibited strong generalizability by providing substantial performance improvements across different speech foundation models. We extract the 15th layer representations from HuBERT and WavLM following EURO [17].
Feature extractor PER on test-clean ($\downarrow$) PER diff.
Initial stage Iter.1-stage1 Iter.1-stage2 after Iter.1
wav2vec 2.0 [3] 19.3% 12.9% 11.9% -7.4%
HuBERT [22] 20.1% 12.8% 10.4% -9.7%
WavLM [11] 18.7% 14.7% 12.7% -6.0%
We evaluate Reborn’s generalizability across different speech foundation models in Table 9. Specifically, in Table 9, we replace the feature extractor from wav2vec 2.0 with HuBERT [22] or WavLM [11]. As presented in the table, Reborn can yield performance improvements regardless of which speech foundation model is used and how well the initial stage performed. The results further strengthen the robustness and the generalizability of our method. Due to the computational resource limitation, we only conduct the ablation for one iteration.
Appendix D wav2vec-U
@@ C.5 Generalizability across Different Speech Foundation Models
wav2vec-U [4] is a popular UASR framework whose variants [17, 28] have achieved SoTA performance on multiple UASR benchmarks. Following prior works [31], wav2vec-U also uses adversarial training to train the phoneme prediction model. The wav2vec-U framework takes the raw waveform as the input and predicts phoneme transcription via three stages: (1) the feature extraction stage, (2) the feature preprocessing stage, and (3) the phoneme prediction stage.
@@
Stage (1) extracts speech features using a self-supervised speech foundation model, wav2vec 2.0 [3]. For a spoken utterance, we denote the extracted feature as $X=[x_{1},x_{2},...,x_{T}]\in\mathbb{R}^{T\times d}$, where $T$ is the time dimension and $d$ is the feature dimension. Stage (2) is the feature preprocessing stage. The dimension of the extracted speech features is first reduced using PCA from $d=1024$ to $d^{\prime}=512$. Next, $k$-means clustering and additional heuristics, namely, adjacent pooling, are adopted sequentially to define the boundary of the segmental structures in the speech features. The features in the same segmental structure are mean-pooled over the two stages into one feature embedding to reduce sequence length from $T$ to $T^{\prime}$ to match phoneme sequences better. In the main text, we call this process "$k$-means-based segmentation". We denote the output from the feature preprocessing stage as $S=[s_{1},s_{2},...,s_{T^{\prime}}]\in\mathbb{R}^{T^{\prime}\times d^{\prime}}$. In Stage (3), the phoneme prediction model takes the pooled speech feature as input and predicts the phoneme sequence $Y=[y_{1},y_{2},...,y_{T^{\prime}}]$. Last, a de-duplication step removes the consecutive duplicate phoneme predictions from $Y$, resulting in a shortened phoneme sequence $Y^{\prime}=[y^{\prime}_{1},y^{\prime}_{2},...,y^{\prime}_{M}]$ without consecutive duplicate phonemes.
During training, the generator $\mathcal{G}$ takes the preprocessed speech feature $S$ as input and generates the logits of $Y$. We can shorten the logits of $Y$ based on its argmax prediction, and the shortened logits (denoted as $G=\mathcal{G}(S)$) are then taken as input by the discriminator $\mathcal{D}$ for GAN training. The adversarial learning loss is defined as follows:
$\mathcal{L}_{GAN}=$ $\min_{\mathcal{G}}\max_{\mathcal{D}}\underset{Z\sim\mathcal{Z}}{\mathbb{E}}[\log(\mathcal{D}(Z))]+\underset{S\sim\mathcal{S}}{\mathbb{E}}[\log(1-\mathcal{D}(\mathcal{G}(S)))]$ (5)
@@
, where $Z$ is a tokenized phoneme sequence sampled from the phonemicized text corpora $\mathcal{Z}$ and $S$ is the preprocessed speech feature sampled from the preprocessed speech corpora $\mathcal{S}$. Besides the min-max objective in the adversarial training, wav2vec-U [4] also adopts the three regularized objectives for better convergence and performance. First, the gradient penalty $\mathcal{L}_{gp}$ [20] is applied on the discriminator $\mathcal{D}$ to stabilize training, where the input of $\mathcal{D}$ is the linear combination of a generated sample $G$ and a real sample $Z$.
$\mathcal{L}_{gp}=\underset{G,Z,\alpha\sim U(0,1)}{\mathbb{E}}\left[\left(\left\|\nabla\mathcal{D}(\alpha G+(1-\alpha)Z)\right\|_{2}-1\right)^{2}\right]$ (6)
@@
Next, the smoothness penalty $\mathcal{L}_{sp}$ is computed across consecutive segments of a generated logit sequence, promoting the generation of closely aligned neighboring outputs by the model.
$\mathcal{L}_{sp}=\sum_{(g_{t},g_{t+1})\in G}\left\|g_{t+1}-g_{t}\right\|^{2}$ (7)
@@
Furthermore, the phoneme diversity loss $\mathcal{L}_{pd}$ is applied to encourage higher vocabulary usage of the generated sequences. As shown in Eq. 8, the entropy of the averaged softmax distribution of the generator outputs ($H_{\mathcal{G}}(\mathcal{G}(S))$) is maximized over the phoneme vocabulary across a batch B.
@@
In this work, we utilize the same training objective from wav2vec-U to perform GAN training (§ 3.2).
Appendix E Implementation Details
E.1 Data Preparation and Phoneme LM Formulation
Our data preparation process mainly follows wav2vec-U [^3] https://github.com/facebookresearch/fairseq/tree/main/examples/wav2vec/unsupervised for fair comparisons. Before the feature extraction stage, we attempt to remove most of the silence through an unsupervised voice activity detection (VAD) method [55]. [^4] https://github.com/zhenghuatan/rVAD However, the method is not completely perfect. To stabilize the GAN training, a special <silence> token is introduced to the text data. The <silence> token is inserted in the front and the end of each sentence, and also between words by a probability of 0.25 according to wav2vec-U. Then, we phonemicize all the words in the text data along with the previously inserted <silence> tokens for GAN training. The two off-the-shelf phonemizers are used to perform phonemicization. For LibriSpeech, we use the G2P phonemeizer [43], and the numerical stress makers are removed to reduce the phoneme set, which is shown to be beneficial for the performance [4, 28, 17]. As for Multilingual LibriSpeech, we use Phonemizer [6], [^5] https://github.com/bootphon/phonemizer a toolkit that supports a wide variety of languages other than English. The language-switching labels are removed as well as the phonemes appear less than 1000 times in the text data. For the Italian, we isolate the double consonant symbol in the lexicon, which is crucial for the initial stage to converge. Note that for TIMIT we do not adopt any of the above data preparations since it originally contains its phone inventories including silence. We use the 39-phone inventory which can be easily found in Kaldi [46] for training and evaluation, [^6] https://github.com/kaldi-asr/kaldi/blob/master/egs/timit/s5/conf/phones.60-48-39.map following prior UASR works [31, 4, 17].
After the data preparation, we directly use the phonemicized text with <silence> token to train the phoneme LM. More specifically, we utilize KenLM [21], [^7] https://github.com/kpu/kenlm to build the 4-gram LM mentioned in 3.1.3. Since in the GAN training stage, we are actually matching the two distributions between the segmented speech features and phonemicized text with <silence>, we naturally build our external LM for RL segmentation learning from the same preprocessed text. The LM is also utilized for the unsupervised cross-validation described in E.5.
E.2 Self-Training with Hidden Markov Models
Originated from semi-supervised learning, self-training aims at providing a trivial approach to utilize the unpaired or unlabeled data effectively [56, 33, 23]. More specifically, given a supervised ASR system $S_{ASR}$ learned from the labeled speech-text data $D_{l}$, we can use $S_{ASR}$ to generate pseudo labels of a large but unlabeled speech data $D_{s}$. Now, we can utilize the speech data along with the pseudo labels to perform supervised learning or re-tuning [56] to obtain a new ASR system $S^{\prime}_{ASR}$. Generally, given that the pseudo label is good enough, we may expect that $S^{\prime}_{ASR}$ performs better than $S_{ASR}$.
In the unsupervised setup, the core concept of self-training is suited as well. By simply using an unsupervised ASR system $S_{UASR}$ to perform pseudo labeling, we can train a new ASR model $S^{\prime}_{UASR}$ by the supervised objective. Since no labeled data is used (we are just using the pseudo labels), the new ASR system does not violate the unsupervised setup. More specifically, we use Hidden Markov Models (HMMs; Bosch et al. [7])as the backbone of the new ASR system. This can be dated back to [10] and is also adopted by recent UASR [4, 28]. Furthermore, to give fair comparisons and increase reproducibility, we follow the publicly available code provided by wav2vec-U for HMM self-training. [^8] https://github.com/facebookresearch/fairseq/tree/main/examples/wav2vec/unsupervised/kaldi_self_train Our results in Table 1 indicate that with the high-quality pseudo labels generated by Reborn, the new ASR model based on HMMs performs better than the original UASR system and is the best among all the other methods.
E.3 Segmentation Model Architecture